The radius of a silver bar is 30 π‘π‘š. What force is required to stretch the bar by


12.7% of its length where Young’s modulus is π‘Œ = 7.25 Γ— 1010 π‘π‘šβˆ’2

1
Expert's answer
2021-12-06T16:42:27-0500

"7.25\cdot10^{10}=\frac{F}{\pi(0.3)^20.127}\\F=2.60\cdot10^{9}\ N"